For each of the following operators T on a vector space V, find an ordered basis B such that [T]e is a diagonal matrix.(a) V = P2 (R) and T(f(x)) = xf(x) + f(2)x+ f(3). db (b) V = M2x2(R) and T b (1 :]) = [.. (c) V = M2x2(R) and T(A) = AT + 2t
Answer #1


@ N=P2 CIR) and ICH(0)) (x) + H2)2 + +(3) Basis Ol Pa CIR ){ 1, 2,222) T(1) = ati (2) - 32+3 7692)278 +49 +9 3 9 [1 4 CA haa v=maxa (IR) b 108 37). Select ** 10.07.Lo!? 20 % [3071 BE matrix A O - Oo - OOO - O eisen values of A) A-dIl=0 - (da) (d+1)=0V= maxa (IR) T(A) = AT tatr(A) Ia Basis 081.[79] [833 [b]- [07-2017007- [:) T T 0 (۵) في 0 [O]• [:]+260°/6°?[60] [!][:7-20(2